Find the unit U(A) of a vector A = -20I + 36J. Given a vector A, U(A) is defined as A/|A|
1
Expert's answer
2020-03-27T14:33:07-0400

Given, A = -20I +36J

"|A|= \sqrt{(-20)^2 + (36)^2}= \sqrt{1696}=4\sqrt{106}"

"U(A) = \frac{A}{|A|} = \frac{-20I+36J}{4\sqrt{106}}=\frac{4(-5I+9J)}{4\sqrt{106}}=\frac{-5I+9J}{\sqrt{106}}"